Fermi National Laboratory


2011 Accelerator Update Archive

December 21, 2011 - December 28, 2011

NuMI reported receiving 8.81E18 protons on target for the period from 12/19/11 to 12/26/11.

  • Muon Rings personnel conducted special high intensity studies
  • New Meson Fermilab Test Beam Facility experiment 1019 began taking beam
  • Booster beam off for short time due to a RF station's transformer suffering a coil short
  • Booster and Main Injector RF problems fixed

December 19, 2011 - December 21, 2011

  • The access for maintenance on 12/20/11 got extended to replace a MI kicker
  • Beam to NuMI was delayed on 12/20/11 due to a Lambertson power supply problem
  • Meson MS2 LCW system had glycol added and was up and running by mid afternoon, 12/21/11

December 16, 2011 - December 19, 2011

  • Linac technician replaced a quadrupole power supply
  • Repairs on a Linac RF station LCW leak held off beam
  • There will be a six-hour access starting at 7 AM on Tuesday, 12/20/11, for maintenance, necessary repairs, and Interlock testing

December 14, 2011 - December 16, 2011

  • FTBF experiment T-1015 taking beam
  • Booster personnel conducted studies
  • OpsCo will have a four-hour access early next week
  • NuMI reported that this was their third best week for protons on target

December 12, 2011 - December 14, 2011

  • Neutron Therapy facility power supply LCW leak repaired
  • A MiniBooNE power supply LCW leak repaired
  • Operators and Booster personnel resolved a problem with console #5 sending interrupts on its own
  • Booster RF station 11 has been tripping off regularly. They will need to access to make rep

December 9, 2011 - December 12, 2011

  • Muon Rings personnel and an Instrumentation person worked on their BPM
  • Meson FTBF experiments T-1015 and T-1017 began taking beam
  • MTA suffers power outage: they weren't taking beam
  • Linac quadrupole power supply problems

December 7, 2011 - December 9, 2011

  • Problems with KRF5.
  • Problems with Booster bulk supply (B:CEP400)

December 5, 2011 - December 7, 2011

  • Booster, Main Injector, and Muon Rings personnel conducted studies
  • Booster kicker personnel repaired a kicker (MSK11) that was holding off beam to Switchyard
  • Operators made adjustments to the I- Source arc current

December 2, 2011 - December 5, 2011

NuMI reported receiving 8.62E18 protons on target for the period from 11/28/11 to 12/5/11.

  • Booster, Main Injector, and Muon Rings personnel conduct studies
  • M01 LCW problems holds off Meson FTBF T-978 experiment for a shot time
  • operators completed a 750 KeB line tune and a Booster injection tune

November 30, 2011 - December 2, 2011

  • A Booster RF station (BRF16) was having trouble
  • MiniBooNE beam turned off during two day shifts for MI-12 service building air conditioning work
  • Booster beam shut off for access and BRF13 repairs

November 28, 2011 - November 30, 2011

  • A string of NuMI Dipole magnets tripped off due to high temperature. Personnel accessed the Main Injector and back flushed the magnets
  • Booster, Main Injector, NuMI, and Muon areas all accessed and conducted maintenance work during Dipole work
  • A Booster RF station (BRF13) tripped off and would not reset. Operators bypassed the station and tuned around it
  • Booster will access on Thursday morning (12/1/11) to repair BRF13
  • MiniBooNE will shut down tomorrow (12/1/11) to install a new air conditioning unit

November 23, 2011 - November 28, 2011

NuMI reported receiving 8.31E18 protons on target for the period from 11/21/11 to 11/28/11.

  • Meson FTBF experiment T-978 pauses their beam while experiment T-1012 take short run over the holidays
  • NuMI BPMs requires reboots
  • Main Injector RF station 15 bypassed
  • Linac RF station 5 down for repairs - halts all beam for ~18 hours
  • Linac RF station 1 down for repairs - halts all beam for ~2 hours

November 21, 2011 - November 23, 2011

  • Linac personnel changed solenoid power supply for a KLystron
  • NuMI personnel accessed their absorber room and cleaned a water skid filter
  • Muon personnel conducted studies
  • MTA took beam

November 18, 2011 - November 21, 2011

NuMI reported receiving 9.22E18 protons on target for the period from 11/14/11 to 11/21/11.

  • Muon Ring (formerly Antiproton Source) studies
  • Problems with a Booster RF station and I- Source vacuum pump trips
  • NuMI suffered with BPM problems
  • This morning, 11/21/11, Linac personnel will work on a Klystron RF station and NuMI will inspect an Absorber room water skid.

November 16, 2011 - November 18, 2011

  • Neutron Therapy Facility treating patients
  • NuMI BPM problems resolved when VME crate failed and was replaced

November 14, 2011 - November 16, 2011

  • A Main Injector beam valve closed due to a compressed air leak and held off beam to FTBF experiment T-978
  • Water Group personnel repaired a Linac Klystron RF station water leak that held off beam for a short time
  • Operators valved out a Linac sensor that had a split water hose, which didn't interrupt beam

November 11, 2011 - November 14, 2011

NuMI reported receiving 7.53E18 protons on target for the period from 11/7/11 to 11/14/11.

  • Muon personnel conducted studies
  • MI Personnel conducted Recycler studies
  • Vacuum problems held off the Main Injector for about 18 hours
  • Lambertson problems held off NuMI for about two hours

November 9, 2011 - November 11, 2011

  • Accesses on Wednesday
  • Booster RF problems
  • T-978 begin taking bean on Wednesday

November 7, 2011 - November 9, 2011

  • RF trips troubled the Booster
  • Problems with the MINOS near detector halted beam at times for NuMI
  • There will be a two hour access this afternoon to fill a kicker with Fluorinert

November 4, 2011 - November 7, 2011

NuMI reported receiving 1.04E19 protons on target for the period from 10/31/11 to 11/7/11. This set a record for being the highest number of protons on target for a week.

  • Neutron Therapy Facility treating patients
  • Booster, and Switchyard personnel conduct studies
  • Operators find and fix liquid nitrogen leak at A0
  • Main Injector gap clearing kicker problems

November 2, 2011 - November 4, 2011

  • Linac, Booster, Main Injector, and Switchyard personnel conduct studies
  • Neutron Therapy Facility treating patients

October 31, 2011 - November 2, 2011

  • Main Injector personnel conducted a Recycler closed orbit study
  • Neutron Therapy Facility personnel treated patients
  • Meson FTBF experiment T-1015 completed its run
  • Main Injector personnel conducting g-2 studies today, 11/2/11

October 28, 2011 - October 31, 2011

  • Neutron Therapy Facility treating patients
  • Antiproton Source personnel conducted Mu2e studies
  • Meson FTBF experiment T-1015 taking beam
  • ComEd power glitch tripped off Linac, Booster, and Main Injector equipment
  • Vacuum pumps trip off in 8 GeV line and
  • Starting on Tuesday, 11/1/11, the Antiproton Source Department will be know as the Muon Department

October 26, 2011 - October 28, 2011

  • Main Injector RF station repairs solved NuMI beam permit problems
  • Meson FTBF experiment T-1015 begins taking beam
  • The Neutron Therapy Facility treats patients
  • Antiproton Source personnel conduct Mu2e studies
  • NuMI sets record for most protons on target in one day, 1.439E18

October 24, 2011 - October 26, 2011

  • MI personnel began g minus 2 (g-2) studies
  • NuMI beam permit trips due to a MI RF station (MIRF17) problem
  • Booster had West Anode Power Supply trips due to BRF9 and 10 problems
  • Antiproton Source personnel conduct reverse proton studies
  • Meson FTBF experiment T-1015 has been cleared to take beam today

October 21, 2011 - October 24, 2011

NuMI reported receiving 8.27E18 protons on target for the period from 10/17/11 to 10/24/11.

  • The Neutron Therapy is treating patients
  • Meson FTBF experiment T-1008 taking beam
  • Main Injector and Antiproton Source personnel conducted studies
  • The Main Injector P1 transfer line suffered a vacuum burst

October 19, 2011 - October 21, 2011

  • Chiller #3 at CUB tripped off, caused the 55 LCW system to warm up, and required Linac to turn off some equipment
  • Two Meson FTBF experiments received permission to take beam: T-1008 and T-1015
  • a Main Injector Micro Wave Absorber had one of its ferrite rods drop into the center of the beam pipe and block beam: Mechanical technicians found and fixe the problem

October 17, 2011 - October 19, 2011

  • A Mechanical Support Engineer lowered the MI-62 LCW system's temperature set limit by five degrees and that seems to have solved the NuMI HV101 magnet trips
  • Pbar personnel conduct proton studies
  • Neutron Therapy Facility treating patients
  • A MI Lambertson magnet power supply continues to cause problems

October 14, 2011 - October 17, 2011

NuMI reported receiving 9.1E18 protons on target for the period from 10/10/11 to 10/17/11.

  • Meson FTBF experiment T-992 taking beam
  • NuMI magnet HV101 high temperature trips under investigation
  • I- Source's sudden current drop repaired by Linac personnel

October 12, 2011 - October 14, 2011

  • Antiproton Source personnel conduct reverse proton studies
  • Booster personnel conducted studies
  • FESS replaces MTA's faulty breaker trip unit
  • MiniBooNE horn power supply's ground fault limit adjust back to nominal

October 10, 2011 - October 12, 2011

  • MiniBooNE off for short time due to Horn problems
  • NuMI off for short times due to ventilation problems
  • MI Lambertson magnet power supply repaired
  • FESS changes out bad MTA breaker trip unit
  • Power glitch tripped off equipment in Linac, Booster, and MI

October 7, 2011 - October 10, 2011

NuMI reported receiving 8.21E18 protons on target for the period from 10/3/11 to 10/10/11.

  • Neutron Therapy Facility treating patients
  • Pbar personnel terminated Pbar's last stack of antiprotons at 11:17 AM on 10/7/11
  • MTA suffers from a power outage
  • MiniBooNE off due to water skid problem and then due to a Horn power supply problem

October 5, 2011 - October 7, 2011

  • NuMI repositioned their target and reconfigured their Horn power supply for antineutrino operations
  • The Pbar LCW leak was discovered
  • Meson FTBF experiment T-922 taking beam
  • Pbar will terminated its stack today, ending the antiproton program at Fermilab

October 3, 2011 - October 5, 2011

  • Antiproton Source personnel conduct studies with their remaining antiprotons
  • Problems with a Switchyard magnet delays beam to FTBF experiment T-992
  • MuMI prepares to move its target to a new position
  • Wednesday (10/5/11) will be a planed maintenance day

September 26, 2011 - October 3, 2011

The Integrated Luminosity for the period from 9/26/11 to 10/3/11 was 49.66 inverse picobarns. NuMI reported receiving 8.21E18 protons on target during this same period.

  • One store provided ~6.5 hours of luminosity to the CDF and D0 experiments
  • NuMI switched their Horn polarity
  • Helen Edwards terminated store 9158 and shut off the Tevatron forever
  • Pbar stopped stacking antiprotons forever
  • A Kautz Rd substation power outage caused the loss of the stash
  • NuMI will prepare to reposition its target on Wednesday (10/5/11) morning

September 28, 2011 - September 30, 2011

  • Four stores provided ~44.75 hours of luminosity
  • Operators rebooted the Main Injector low level RF system
  • The next Meson FTBF experiment T-1012 began taking beam

September 30, 2011

  • Special Edition of the Operation Summary

September 26, 2011 - September 28, 2011

  • Three stores provided ~46.5 hours of luminosity
  • Linac has Klystron vernier problems
  • Pelletron tripped off, but reset
  • Tevatron Electron Lens study
  • MiniBooNE surpassed 1E21 Protons On Target in antineutrino mode

September 23, 2011 - September 26, 2011

The Integrated Luminosity for the period from 9/19/11 to 9/26/11 was 58.32 inverse picobarns. NuMI reported receiving 2.1E18 protons on target during this same period.

  • Five stores provided ~62 hours of luminosity
  • Meson FTBF experiment T-1017 ended their run
  • Access needed to fill Recycler kicker with Fluorinert
  • NuMI resumed taking beam on Saturday morning (9/24/11)
  • FESS personnel installed a portable air conditioner at sector C1 Electronics
  • Operators performed MI coalescing checks; everything looked good

September 21, 2011 - September 23, 2011

  • Three stores provided ~40.25 hours of luminosity
  • A new Meson FTBF experiment T-1017 began taking beam
  • MI and Instrumentation personnel tested new BPM software
  • NuMI personnel completed a low intensity scan of their newly installed target
  • Tevatron personnel conducted a study using two high intensity proton bunches
  • NuMI personnel reported leaks in their target cooling system that need repair. This will delay their taking beam until Saturday, 9/234/11

September 19, 2011 - September 21, 2011

  • Two stores provide ~14 hours of luminosity
  • Cryo System personnel repair the Tevatron sector F3 wet engine
  • Meson FTBF Meson experiment T-922 ended their run
  • The Tevatron quenched during a shot setup on a separator spark

September 16, 2011 - September 19, 2011

The Integrated Luminosity for the period from 9/12/11 to 9/19/11 was 27.36 inverse picobarns. NuMI reported receiving 2.6E18 protons on target during this same period.

  • Five stores provided ~43.75 hours of luminosity
  • Tevatron completed its run of 450 GeV 3x3 study stores
  • Pbar Debuncher kicker pulsing problem fixed
  • The MiniBooNE Horn tripped off due to an strip line air pressure sensor
  • Store 9131 aborted due to a problem with the C3 dump
  • Store 9135 quenched due to a kicker prefire

September 14, 2011 - September 16, 2011

  • The Tevatron completed its run of 450 GeV 3x3 study stores
  • The Tevatron quenched twice during wet squeezes
  • Booster RF suffered from a short communications problem
  • The next Meson FTBF experiment, T-922, began taking beam
  • A Pbar Debuncher kicker had its power supply replaced

September 12, 2011 - September 14, 2011

  • One store provided ~ 5 hours of luminosity
  • Tevatron personnel conduct two more 450 GeV 3x3 study stores
  • MI ground fault finally found and repaired
  • Tevatron cryo system down during wet engine repair
  • Power glitch caused problems with Linac RF, Booster RF, Main Injector kickers, and the Tevatron ramp
  • NuMI personnel will begin replacing their target tomorrow morning, 9/15/11

September 9, 2011 - September 12, 2011

The Integrated Luminosity for the period from 9/5/11 to 9/12/11 was estimated at 47.62 inverse picobarns. NuMI reported receiving 8.74E18 protons on target during this same period.

  • Four stores provided ~28 hours of luminosity
  • MI accessed to fill Kicker with Fluorinert
  • Tevatron personnel conducted many 450 GeV low energy study stores
  • NuMI suffered from beam permit drops
  • Store 9097 aborted due to a bad CAMAC card
  • Store 9099 aborted due to a standby power unit failure

September 7, 2011 - September 9, 2011

  • Two stores provided ~25 hours of luminosity
  • Low energy Tevatron studies
  • PA for LRF1 was replaced

September 5, 2011 - September 7, 2011

  • Three stores provided ~36.75 hours of luminosity
  • A Booster RF station (BRF7) tripped off many times
  • The Tevatron quenched during shot setup due to a kicker prefire
  • MI RF station (MIRF3) trips caused many NuMI beam permit trips
  • Controls personnel repaired problems with the MI-30 Beam Position Monitors

September 2, 2011 - September 5, 2011

  • Five stores provided ~62 hours of luminosity
  • Linac I- Source vacuum valve closed
  • Booster RF system trips due to bad timing card
  • Tevatron quench during shot setup due to kicker prefire
  • Pbar tripped off due to a target blower motor failure

August 31, 2011 - September 2, 2011

  • Three stores provided ~33 hours of luminosity
  • Pbar kicker cables re-terminated (D:EKIK)
  • Tevatron suffered a beam abort due to bad vacuum and a quench due to kicker problems
  • Tevatron RF cavity (TRF5) LCW leak required access to sector F0 to clean up and repair
  • Fermilab lowers power usage for ComEd power test

August 26, 2011 - August 29, 2011

  • Eight stores provided ~58.5 hours of luminosity
  • Booster RF work completed
  • Linac RF station #3 repaired
  • MiniBooNE Horn problem fixed
  • Pelletron suffered a full discharge
  • Antiproton injection to Tevatron lost due to kicker prefire

August 24, 2011 - August 26, 2011

  • Two stores provided ~25.75 hours of luminosity
  • Pelletron problems continue
  • Booster used access to repair two RF stations
  • Two Tevatron 3x3 study stores ended with a quench and an abort
  • MiniBooNE horn charging power supply problem delayed beam
  • Run Coordinator will allow Linac and Booster maintenance work today

August 22, 2011 - August 24, 2011

  • One store provided ~8.5 hours of luminosity
  • Tevatron Transfer Hall vacuum leak repaired
  • Tevatron personnel conduct 3x3 studies
  • Antiproton source target air blower repaired
  • Booster West Anode Power Supply problems
  • Pelletron solenoid problems

August 19, 2011 - August 22, 2011

The Integrated Luminosity for the period from 8/15/11 to 8/22/11 was 31.61 inverse picobarns. NuMI reported receiving 6.63E18 protons on target during this same period.

  • Five stores provide ~29.75 hours of luminosity
  • Booster had RF and kicker problems
  • D0 cold compressor tripped off
  • Tevatron aborted once and quenched once
  • Tevatron developed A17 vacuum leak
  • Pelletron back on line

August 17, 2011 - August 19, 2011

  • Four stores provided ~23.25 hours of luminosity
  • Store 8996 was a 3X3 study store and lost due to a quench
  • Pelletron suffered full discharge and put in recovery mode
  • ComEd offsite power glitch caused many devices to trip off

August 15, 2011 - August 17, 2011

  • Two stores provided ~20.5 hours of luminosity and two study stores that did not provide any useful luminosity
  • Tevatron Kautzky valve repair held off beam
  • Pelletron conditioning held off Recycler stochastic cooling
  • Problematic ACNET code halted transfer to Recycler

August 12, 2011 - August 15, 2011

  • Two store provided ~47.75 hours of luminosity
  • Massive water leak in the Linac gallery
  • Two Tevatron quenches

August 10, 2011 - August 12, 2011

  • Five stores provided ~37.75 hours of luminosity
  • Ground water leak found in the 750 KeV area
  • Pelletron brought back online
  • Tevatron Collimator issue resolved
  • Tevatron suck at 90 GeV due to bad UPS in sector A2

August 8, 2011 - August 10, 2011

  • Six store provided ~ 41.5 hours of luminosity
  • Pelletron problems continue
  • Many NuMI beam permit trips due to Main Injector RF problems, but all the RF stations were on and operating
  • Resistivity of the Antiproton Source 95 Degree LCW system was dropping; FESS change deionizing bottles

August 5, 2011 - August 8, 2011

The Integrated Luminosity for the period from 8/1/11 to 8/8/11 was 5.82 inverse picobarns. During the same time period NuMI took 6.57E18 of beam.

  • Five stores provided ~20.25 hours of luminosity
  • Tevatron cryo systems and vacuum ready to go
  • Tevatron quenched during turn on
  • Store 8954 aborted due to loss of power to control rack
  • Pelletron problems continue to delay its use
  • MTA hopes to take beam this morning (8/8/11)

August 3, 2011 - August 5, 2011

  • Antiproton Source LCW leak found and fixed
  • New Meson Fixed Target Test Beam experiment T-1010 begins taking beam
  • MTA taking beam
  • ComEd power glitch tripped off Linac, Booster, and Main Injector; everything reset
  • Work continues on the Recycler Pelletron
  • Tevatron should be ready to establish a store later this evening

August 1, 2011 - August 3, 2011

  • Tevatron still in cryo system and vacuum recovery mode
  • Much maintenance work being done throughout the accelerator complex
  • Pbar resumes stacking
  • Lightning strike ends the Meson FTBF T-1008's current run at Fermilab
  • Cryo Coordinator hopes the Tevatron cryo system will be ready to go on Friday (8/5/11)

July 29, 2011 - August 1, 2011

The Integrated Luminosity for the period from 7/25/11 to 8/1/11 was 17.71 inverse picobarns. During the same time period NuMI took 8.8E17 of beam.

  • There were no stores due to the Tevatron cryo systems being in recovery mode
  • Main Substation brought back online
  • Operators and accelerator personnel work to recover all system
  • Beam established to NuMI, MiniBooNE, the Meson FTBF T-1008 experiments
  • Controls problems halted MTA from taking beam

July 27, 2011 - July 29, 2011

  • One store provided ~12.75 hours of luminosity
  • NuMI magnet cabling work continues
  • The Main Accelerator Safety System dropped
  • The Master Substation and Kautz Road Substation lost power due to ComEd power lines going down in storm
  • Many accelerator devices had to be replaced or repaired
  • Tevatron cryo and vacuum systems recovery will take three to six days

July 25, 2011 - July 27, 2011

  • Three stores provided 31.5 hours of luminosity
  • Pbar Kicker personnel repaired abort kicker
  • Store 8927 quenched due to CDF beam abort
  • Main Injector and NuMI down for (V108) cable pulls
  • Linac personnel repair RF station (LRF3)
  • Pbar accessed to clean clogged fan exhaust on Debuncher RF cavity (DRF3)

July 22, 2011 - July 25, 2011

  • Four stores provide ~28.2 hours of luminosity
  • Lightning strikes caused the loss of stores 8912 and 8923
  • NuMI magnet V108 has six high voltage cables, out of sixteen, grounded together and must be replaced
  • Lightning tripping off equipment and systems everywhere
  • Hoses for MiniBooNE quadrupole magnet (Q847) found clogged

July 20, 2011 - July 22, 2011

  • Three stores provided ~39.25 hours of luminosity
  • Pbar LCW leak found and fixed
  • MI LCW leak required LCW pumps turned off
  • MI LCW leak found and fixed
  • NuMI V108 magnet/power supply cabling problem

July 18, 2011 - July 20, 2011

  • Three stores provided ~29.25 hours of luminosity
  • High temperatures and failing air conditioners keep FESS personnel on the go
  • High temperatures have caused many equipment and beam problems
  • Cryo system technicians repaired the Tevatron sector D1 wet engine flywheel bearings
  • Store 8904 aborted due to Tevatron quadrupole trip
  • Booster personnel repaired the 400 MeV chopper
  • New Meson FTBF T-1008 experiment begins taking beam
  • NuMI hopes to resume taking normal beam on Thursday, 7/21/11

July 15, 2011 - July 18, 2011

The Integrated Luminosity for the period from 7/11/11 to 7/18/11 was 49.33 inverse picobarns. NuMI is off for target replacement; they expect to resume taking beam this Friday, 7/22/11.

  • Four stores provided ~48.25 hours of luminosity
  • High temperatures in MiniBooNE's MI-12 service building require power supplies to be shut off
  • MI coalescing studies
  • Regeneration of CUB's column #1 halted due to water leak
  • Pelletron suffered many trips
  • Air Conditioning problems occurred at many places
  • Main Injector/Tevatron injection problems
  • The plans for today are to survive the heat

July 13, 2011 - July 15, 2011

  • Four stores provided ~21 hours of luminosity
  • Tevatron heat exchanger work continues
  • MI LCW leak found and fixed
  • Pbar LCW leak still a mystery
  • Store 8891 aborted due to Tevatron vacuum valves
  • MTA took beam
  • Search for Tevatron ground fault halted when fault disappeared
  • KRF6 and LRF4 problems held off shot setup for approximately six hours

July 11, 2011 - July 13, 2011

  • Two stores provided ~15.25 hours of luminosity
  • Lightning strikes cause equipment problems
  • Water technicians work on Tevatron heat exchangers
  • MI damper problem fixed its self - probably a loose connector
  • Store 8880 intentionally aborted due to Tevatron E3 wet engine loosing its belt
  • E3 wet engine needed complete overhaul
  • MI-12 service building still suffers from temperature problems
  • MTA conducts first test of beam through its RF cell filled with nitrogen
  • EE Support personnel work on the MI Lambertson power supply
  • Meson MTest experiment T-1015 completes run at Fermilab
  • Run Coordinator will allow MI an access to search for LCW leak today (7/13/11)

July 8, 2011 - July 11, 2011

The Integrated Luminosity for the period from 7/4/11 to 7/11/11 was 58.1 inverse picobarns. NuMI reported receiving 3.55E18 protons on target during this same period.

  • Four stores provided ~33.25 hours of luminosity
  • NuMI conducted a target scan and found water in their target. NuMI is off for about 2.5 weeks for a target change
  • Since the Tevatron needed vacuum work, much of the accelerator complex conducted maintenance during the day shift on 7/8/11
  • MI suffered from coalescing and damper problems
  • MiniBooNE off due to high temperatures in the MI-12 service building
  • Stash lost on 7/11/11 due to lightning strike
  • Run Coordinator allows Tevatron heat exchanger work and other jobs during the day shift that would not hinder stacking and stashing

July 6, 2011 - July 8, 2011

  • Three stores provided ~30.75 hours of luminosity
  • MI kicker (I:KPS5S) repaired
  • Relay replaced for Tevatron sector D3 LCW pump
  • Pbar safety system testing completed
  • Tevatron kicker E48 problems began
  • FESS reported problems with Tevatron sector C3 pond water pump
  • Stash lost due to MI-60 LCW pumps tripping off
  • Today (7/8/11), Run Coordinator will allow access for Tevatron vacuum work and CDF and D0 access, for Tevatron kicker work, for a NuMI target scan, for Booster studies, and for a thermal scan of the MI-60 LCW pump motors and breaker panels

July 4, 2011 - July 6, 2011

  • Three stores provided ~40 hours of luminosity
  • MI-12 service building air conditioner tripped off, but MiniBooNE continued taking beam while the building cooled down
  • Cryo system technicians tighten belts on Tevatron wet engines
  • Main Injector, Recycler, Pelletron, Booster, and MTA all accessed for various maintenance work
  • Operators investigating high LCW temperatures at C1 & 2 found three-way valve set in wrong position
  • MI kicker's Thyratron failed
  • Run Coordinator shuts off Tevatron on 6/6/11 for D0 and CDF access, Pbar and Tevatron safety system testing, and for the 6-8 hour replacement of a MI kicker's Thyratron

July 1, 2011 - July 4, 2011

The Integrated Luminosity for the period from 6/27/11 to 7/4/11 was 59.43 inverse picobarns. NuMI reported receiving 5.12E18 protons on target during this same period.

  • Five stores provided ~ 64.5 hours of luminosity
  • Tevatron sector E48 kicker causing problems
  • NuMI dehumidifier off due to problems for weekend
  • Power glitch tripped off many devises
  • MI-12 service building high temperature caused MiniBooNE to turn off for seven hours
  • A1 transfer line suffered a vacuum burst
  • Linac RF station (LRF2) problems held off beam for an hour

June 29, 2011 - July 1, 2011

  • Three stores provided ~29 hours of luminosity
  • Safety system testing continues throughout the accelerator complex
  • Linac personnel repair LRF4 & 5 and the I- Source
  • Tevatron quenched during injection of antiprotons due to air compressor trip that caused kicker misfire
  • New Meson FTFB experiment T-1015 begins taking beam

June 27, 2011 - June 29, 2011

  • Three stores provided ~33.25 hours of luminosity
  • Safety system testing site wide
  • BRF16 sparking problems
  • Switchyard vacuum problems
  • NuMI target scans
  • Tevatron cryo system wet engine work
  • Meson FTBF experiment T-978 completes run
  • Minor power glitch trips off power supplies and other devices
  • Booster off for RF work on 6/29/11

June 24, 2011 - June 27, 2011

The Integrated Luminosity for the period from 6/20/11 to 6/27/11 was 66.87 inverse picobarns. NuMI reported receiving 4.35E18 protons on target during this same period.

  • Four stores provided ~ 59.07 hours of luminosity
  • Pbar Vault exhaust fan repaired
  • Store 8836 aborted due to Tevatron LCW problem
  • MTA took beam
  • I- Source tripped off due to water resister water skid trip
  • Linac off due to LRF3 & 4 power amplifier troubles
  • RF personnel repair BRF12
  • Tevatron abort due to Heater Firing Unit failure
  • Linac off due to door interlock fault, followed by Linac safety system testing

June 22, 2011 - June 24, 2011

  • Three stores provided ~ 34 hours of luminosity
  • NuMI power supply E:V118 lower trip limit lowered
  • CDF and D0 access their collision halls
  • Tevatron RF station (TRF4) had its LCW filter changed
  • Beam aborted during shot setup
  • MTA (MuCool) took an hour of beam
  • A loose cable connection on BFR5 causes Booster problems
  • Today: FESS will repair Pbar exhaust fan, Booster will work on BRF5, MTA will test a power supply, and NuMI will conduct a target scan

June 20, 2011 - June 22, 2011

  • Four stores provided 44.25 hours of luminosity
  • NuMI personnel conducted target scans
  • Booster beam tripped off when an RF equipment rack lost its power
  • Proton Source beam suffered from Booster West Anode power supply trips
  • Interlock technicians begin conducting safety system tests of the accelerators
  • Tevatron personnel conduct luminosity leveling studies
  • NuMI reports possibly finding the cause of their helium leak

June 17, 2011 - June 20, 2011

The Integrated Luminosity for the period from 6/13/11 to 6/20/11 was 78.44 inverse picobarns. NuMI reported receiving 5.64E18 protons on target during this same period.

  • Five Stores provided ~62.5 hours of luminosity
  • Cryo system technicians repair the Tevatron sector F4 wet engine flywheel
  • Pbar and RF personnel resolve air-flow cooling problem for Accumulator RF station
  • Booster RF station BRF6 bypassed and replace with BRF19
  • Two transfers of antiprotons lost
  • Booster extraction kicker problem repaired
  • NuMI beam tripped off when helium flow through the target increased - it was reset
  • NuMI personnel conducting target scans today (6/20/11). They will move the target on Tuesday (6/21/11)

June 15, 2011 - June 17, 2011

  • The MI phasing controller still causing problems
  • Tevatron Electron Lens (TEL) Front End work completed
  • MiniBooNE conducts air pressure test of MI-12 service building
  • Meson water-to-air cooling tower problems
  • Tevatron personnel conduct hollow beam collimation
  • Pbar Debuncher Lens LCW flow meter problem discovered

June 13, 2011 - June 15, 2011

  • Four stores provided ~37.25 hours of luminosity
  • Cryo system technicians worked on the Tevatron sector B2 system
  • The Main Injector Excitation Control And Regulation system causes problems
  • Store 8816 set second highest luminosity mark with 424.56E30
  • Tevatron Hollow Beam studies
  • Store 8816 lost due to lightning strike

June 10, 2011 - June 13, 2011

  • Six stores provided ~62 hours of luminosity
  • Store 8805 aborted due to a power supply trip
  • NuMI began receiving beam on Saturday morning (6/11/11)
  • Store 8809 aborted due to a vacuum indication
  • Booster beam instability fixed

June 8, 2011 - June 10, 2011

  • Three stores provided ~15 hours of luminosity
  • FESS personnel switched to different power cables for the Tevatron sector A0 main breaker
  • Store 8797 lost due to Tevatron quench
  • Store 8799 aborted due to lightning strike
  • Store 8801 aborted due to lightning strike
  • Tevatron and Main Injector vacuum problems fixed
  • Heavy rains caused flooding problems
  • Tevatron personnel conducted Electron Cloud studies
  • Feeder #36 fault caused power loss for part of Proton area
  • NuMI should resume taking beam on Saturday, 6/11/11

June 6, 2011 - June 8, 2011

  • Three stores provided ~44.65 hours of luminosity
  • Mechanical support personnel installed a portable air compressor in the Main Injector
  • Booster personnel removed NorthStar experimental equipment from MI-8
  • 123E10 of antiprotons lost due to MI vacuum burst
  • Store 8797 aborted due to power cable fault at Tevatron sector A0
  • NuMI hopes to be taking beam by 6/11/11

June 3, 2011 - June 6, 2011

  • Five stores provided ~67 hours of luminosity
  • Kicker trip during transfer caused loss of antiprotons
  • Booster and MiniBooNE have problems due to high temperatures that reduced their intensities
  • MTA taking beam
  • The Booster NorthStar experiment ended on Sunday morning

June 1, 2011 - June 3, 2011

  • Three stores provided ~22.75 hours of luminosity
  • NorthStar isotope experiment installed in MI-8 line for a three day run
  • Cryo system technicians overhauled two Tevatron wet engines
  • Store 8787 quenched due to a computer problem that moved collimators into the colliding beams
  • Linac chopper repaired

May 30, 2011 - June 1, 2011

  • Three stores provided ~40.75 hours of luminosity
  • MI-52 Kicker problem fixed
  • Electron Cooling problems slowed shot setup for store 8782
  • Lightning strike caused loss of store 8782 - No quench
  • MI RF troubled by vacuum burst
  • MTA (MuCool) took beam
  • Tevatron access today (6/1/11) for wet engine repairs

May 27, 2011 - May 31, 2011

The Integrated Luminosity for the period from 5/23/11 to 5/30/11 was 58.41 inverse picobarns.

  • Five stores provided ~41.5 hours of luminosity
  • NuMI off for target replacement
  • Tevatron Electron Lens noise caused beam problems
  • Tevatron sector B3 wet engine repairs halted - B3 cryo system operated via Jewel Thompson (JT) valve
  • Store 8770 quenched due to abort kicker prefire
  • Lightening strikes interrupts shot setup and disrupts other systems

May 25, 2011 - May 27, 2011

  • Four stores provided ~37.5 hours of luminosity
  • Tevatron personnel conduct crystal collimator studies
  • Jasmin experimenters finish studies in the Antiproton Source
  • NuMI personnel complete Horn scans
  • D0 dry engine repaired

May 23, 2011 - May 25, 2011

  • Four stores provided ~23.5 hours of luminosity
  • Tevatron personnel conduct crystal collimator studies
  • Recycler's Electron Cooling communications restored

May 20, 2011 - May 23, 2011

The Integrated Luminosity for the period from 5/16/11 to 5/23/11 was 59.22 inverse picobarns. NuMI took 4.73E16 of beam during this same time period.

  • Four stores provide ~47.25 hours of luminosity
  • NuMI off for target repairs
  • Tevatron personnel conduct hollow beam studies
  • Tevatron quench during transfer of antiprotons
  • EE Support repairs Tevatron quadrupole regulator's load cable
  • Lightening strike on the evening of May 22, caused Recycler kickers to prefire, the loss of the stash, and lost communication with Pelletron

May 18, 2011 - May 20, 2011

  • Four stores provide ~ 37.5 hours of luminosity
  • Tevatron personnel conduct several crystal collimator studies
  • Tevatron personnel conduct a hollow beam study
  • Operations reported network storms
  • Run Coordinator allows CDF a three-hour access this morning, 5/20/11

May 16, 2011 - May 18, 2011

  • Three stores provided ~14.25 hours of luminosity
  • Tevatron wet engines at A3 and F3 repaired
  • NuMI's target failed. Experiment off for one month
  • Tevatron D2 wet engine blows seal, caused store 8741 to quench
  • Tevatron quenched during checkout when Quench Protection Monitor rebooted itself
  • Tevatron personnel conducts crystal collimator study

May 13, 2011 - May 16, 2011

The Integrated Luminosity for the period from 5/9/11 to 5/16/11 was 63.88 inverse picobarns. NuMI took 4.15E18 of beam during this same time period.

  • Five stores provided ~63 hours of luminosity
  • Tevatron personnel conduct Hollow Beam studies
  • EE Support personnel help Operations turn on the Tevatron Electron Lens
  • Booster suffered from low-level RF room temperature problems
  • NuMI target problems held off NuMI beam
  • Necessary cryo system repairs on wet engine flywheels required the Run Coordinator to terminated store 8737 this morning, May 16th

May 11, 2011 - May 13, 2011

  • Three stores provided ~19.25 hours of luminosity
  • FTBF experiment T-978 completed their run on May 11th
  • Power glitch affected many systems
  • Lightning strike caused partial loss of stash and tripped off equipment
  • MI quadrupole tripped off on ground fault indication - bus problem repaired by Mechanical Support
  • Pbar personnel replace the target
  • FTBF experiment T-979 began their run on May 13th

May 9, 2011 - May 11, 2011

  • Four stores provided ~37.5 hours of luminosity
  • Network storms and Linac RF problems delayed termination of store 8719
  • Pbar sweeper magnet controls problems halted stacking for an hour
  • A Switchyard power supply failure held off beam to FTBF experiment T-978 for about three hours

May 6, 2011 - May 9, 2011

The Integrated Luminosity for the period from 5/2/11 to 5/9/11 was 76.47 inverse picobarns. NuMI took 3.98E18 of beam during this same time period.

  • Three stores provided ~63.25 hours of luminosity
  • Linac personnel repaired multiple problems with LRF1 & 5
  • Network storms delayed shot setup

May 4, 2011 - May 6, 2011

  • Three stores provide ~39.75 hours of luminosity
  • Tevatron personnel conduct hollow beam studies
  • Linac personnel repairs KRF7's LCW leak
  • Operators tuned to correct MI-8 collimator beam positions
  • The Thursday day shift was full of accesses for repairs
  • MIRF7 tripped off and operators bypassed the station
  • Cryo system personnel replaced a wet engine flywheel at F2 on Friday, 5/6/11

May 2, 2011 - May 4, 2011

  • Three stores provided ~45.5 hours of luminosity
  • Controls personnel repair Pbar repeater problem
  • Store 8709 sets record with a luminosity of 431.07E30

April 29, 2011 - May 2, 2011

The Integrated Luminosity for the period from 4/25/11 to 5/2/11 was 71.36 inverse picobarns. NuMI took 4.06E18 of beam during this same time period.

  • Four stores provided ~64 hours of luminosity
  • Store 8701 set a new record with a initial luminosity of 416.61E30
  • Booster RF system trips caused some minor downtime
  • Pbar Debuncher kicker repairs interrupted stacking for three hours
  • A beam instability at the beginning of store 8702 cause a loss of luminosity

April 27, 2011 - April 29, 2011

  • Three stores provided ~43.25 hours of luminosity
  • Tevatron RF station 4 repaired in-between stores
  • Operators tuned on the I- Source output
  • Store 8700 set a new record with a luminosity of 405.25E30
  • Operators monitored the conditioning for a New Muon Lab coupler cavity

April 22, 2011 - April 25, 2011

The Integrated Luminosity for the period from 4/18/11 to 4/25/11 was 49.01 inverse picobarns. NuMI took 3.36E18 of beam during this same time period.

  • Four stores provided ~41.25 hours of luminosity
  • Cryo system personnel replaced a fly wheel and bearings on two Tevatron wet engines
  • Linac personnel repaired a loose drive filament connection on LRF2
  • D-Zero mad several accesses for necessary repairs
  • Linac personnel repaired a vacuum leak on RF tank #2
  • Pbar worked on an Accumulator emittance problem
  • The Tevatron quenched with no store in the machine

April 20, 2011 - April 22, 2011

  • Three stores provided ~34 hours of luminosity
  • MI-20 LCW pump brought back online
  • MuCool studies
  • MuCool took high intensity beam to their target
  • Linac RF station power amplifier replaced
  • Cryo system personnel will replace the flywheel for Tevatron sector A4 wet engine at 10 AM, Friday, 4/22/11

April 18, 2011 - April 20, 2011

  • Three stores provided ~22.75 hours of luminosity
  • NuMI beam intensity raised to 1.6E13 protons per pulse
  • Store 8666 quenched due to a control card failure
  • A lightning strike aborted store 8672
  • Tevatron and MuCool studies

April 15, 2011 - April 18, 2011

The Integrated Luminosity for the period from 4/11/11 to 4/18/11 was 36.6 inverse picobarns. NuMI took 4.9E18 of beam during this same time period.

  • Five stores provided ~61.5 hours of luminosity
  • Pbar personnel work on their dump LCW system
  • NuMI beam intensity lowered to 1.4E13 due to target LCW leak
  • MiniBooNE beam intensity lowered due to MiniBooNE and MI beam permit trips

April 13, 2011 - April 15, 2011

  • There was no delivered luminosity for this period due to Tevatron kicker work
  • Linac and Booster personnel conduct Neutron Therapy Facility beam studies
  • Tevatron kicker work completed on Thursday night (4/14/11)
  • MuCool experiment took some beam

April 11, 2011 - April 13, 2011

  • Two stores provided ~16.75 hours of luminosity
  • MI personnel brought machine back from RF system trip
  • Tevatron quench during shot setup due to kicker prefire
  • NuMI beam permit trips due to Lambertson power supply high temperatures
  • MI LCW tripped off and caused the loss of the Recycler stash
  • Store 8641 quenched due to kicker prefire
  • Tevatron kickers put in conditioning mode
  • Kicker personnel repaired Recycler kicker

April 8, 2011 - April 11, 2011

  • Five stores provided ~66.5 hours of luminosity
  • NuMI personnel conduct target scans
  • MI personnel conduct G minus 2 studies
  • Pbar dump LCW system developed leak
  • NuMI begins taking beam
  • Linac Klystron RF station's charging power supply repaired
  • Linac RF station repaired

April 6, 2011 - April 8, 2011

  • Three stores provided ~34.75 hours of luminosity
  • Main Injector ramp trip caused the transfer loss of antiprotons to the Recycler
  • CALICE experiment T-978 returned to the Fermilab Test Beam Facility
  • NuMI conducting target scans, will resume normal operations on Saturday, 4/9/11

April 4, 2011 - April 6, 2011

  • Two stores provide ~31.25 hours of luminosity
  • CDF and D0 accessed their collision halls
  • FTFB T-932 experiment completed their run
  • Tevatron access to investigated sump pump and vacuum pump problems

March 30, 2011 - April 4, 2011

The Integrated Luminosity for the period from 3/28/11 to 4/4/11 was 36.95 inverse picobarns. NuMI was off to replace its target.

  • Seven stores provided ~83.25 hours of luminosity
  • Tevatron quench during shot setup due to RF station trip
  • Store 8606 quenched due to vertical corrector magnet power supply glitch
  • Proton Source beam quality problems being investigated
  • New Meson FTFB experiment T-932
  • Tevatron quench during shot setup due to RF station trip
  • Store 8603 quenched due to vertical corrector magnet glitch
  • Proton Source investigating beam position problems
  • New Meson FTFB T-932 experiment began taking beam
  • MI LCW magnet leak found and fixed
  • Tevatron quench caused by lightening strike

March 28, 2011 - March 30, 2011

  • No stores due to D16-2 magnet replacement
  • D16-2 magnet replacement is complete
  • Tev is cold
  • Main Ring, Booster and F-Sector acceses

March 25, 2011 - March 28, 2011

  • No stores due to D16-2 magnet replacement
  • Booster and Main Injector RF Problems

March 23, 2011 - March 25, 2011

  • No stores due to D16-2 magnet replacement
  • Main Injector water leak
  • Booster, Main Injector, and F-Sector accesses

March 21, 2011 - March 23, 2011

  • No stores due to D16-2 magnet replacement
  • New Recycler stash record (608.6E10)
  • RF problems in Linac and Main Injector

March 18, 2011 - March 21, 2011

The Integrated Luminosity for the period from 3/14/11 to 3/21/11 was 70.75 inverse picobarns. NuMI is off for target replacement.

  • Three stores provided ~44 hours of luminosity
  • Time Line Generator problems fixed
  • Tevatron RF station trip caused luminosity loss
  • Controls personnel replaced bad Meson Septa motors
  • Booster personnel repaired chopper
  • Lightning strikes caused store 8594 to abort and many other accelerator problems
  • Tevatron cryo magnet failed and must be replaced - Tevatron off for 10-11 days

March 16, 2011 - March 18, 2011

  • Four stores provide ~44.25 hours of luminosity
  • Recycler kicker problems
  • Operators filling Antiproton Source water system
  • FESS configured the MI-30 LCW pump to run remotely

March 14, 2011 - March 16, 2011

  • Fours stores provided ~39.31 hours of luminosity
  • Recycler kicker pre-fired
  • MiniBooNE resumed taking beam 3/15/11 during day shift
  • Meson Septa motors replaced

March 11, 2011 - March 14, 2011

The Integrated Luminosity for the period from 3/7/11 to 3/14/11 was 6.18 inverse picobarns. NuMI is off for target replacement.

  • Three stores provided ~20.75 hours of luminosity
  • Main Injector repaired transductor problem
  • Antiproton Source power supply replaced
  • Tevatron repairs higher order power supply problem
  • Instrumentation technician repairs Tevatron BPM Front End
  • Store 8571 lost due to a quench
  • Tevatron sector B0 cryo system compressor failed

March 9, 2011 - March 11, 2011

  • The accelerator shut down, due to ComEd work, continued
  • Personnel at the Central Utility Building rotated LCW pumps for maintenance work
  • Water Work personnel found and fixed a Main Injector leak
  • Operators monitored the New Muon Lab's cryo cavity coupler conditioning
  • A cryo compressor in Tevatron F-sector tripped off
  • Operators monitored the conditioning of the Tevatron separators
  • Operations reported that the TEV magnet Tilt Meters detected the magnitude 8.9 earthquake off the coast of Japan

March 7, 2011 - March 9, 2011

  • Due to ComEd work, Fermi accelerators are in shutdown
  • Tevatron, Operators, and other personnel collect Tevatron magnet roll measurements
  • FESS personnel replace breaker at Tevatron sector D3
  • Central Utility Building personnel conduct pump maintenance

March 4, 2011 - March 7, 2011

The Integrated Luminosity for the period from 2/28/11 to 3/7/11 was 52.1 inverse picobarns. NuMI was off to replace its target.

  • Four stores provided ~41 hours of luminosity
  • Store 8549 quenched due to a kicker problem
  • A Switchyard Septa tripped off due to a vacuum problem
  • The Tevatron quenched during shot setup due to another kicker problem
  • Linac and Operator personnel tuned up the I- Source
  • Tevatron and Antiproton Source personnel conducted studies
  • Due to ComEd work, Operations turned off the accelerator complex for machine upgrades and repairs

March 2, 2011 - March 4, 2011

Four stores provided ~34.5 hours of luminosity

  • Store 8542 quenched due to a kicker pre-fire
  • A new Meson MTest experiment, T-1012 (a TAU neutrino Test) began
  • The accelerator complex will be down from 3/7/11 to 3/11/11 due to ComED work
  • The machines will perform upgrades and necessary repairs during this downtime

February 28, 2011 - March 2, 2011

  • Three stores provided ~42.5 hours of luminosity
  • Two Main Injector electrical safety system ground faults caused the loss of two stashes
  • Tevatron personnel will conduct an end-of-store study on store 8542

February 25, 2011 - February 28, 2011

The Integrated Luminosity for the period from 2/21/11 to 2/28/11 was 41.81 inverse picobarns. NuMI reported receiving 3.98E18 protons on target during this same period.

  • Four stores provided ~33.35 hours of luminosity
  • NuMI target cooling leak investigation continues
  • Tevatron personnel conducted end-of-store studies
  • Linac personnel repairs low-energy Buncher
  • MI resolves power supply phasing problems
  • D0 required eight-hour access for repairs
  • Store 8543 lost due to lightning strike
  • Shot setup halted due to RF trip and quench

February 23, 2011 - February 25, 2011

  • Two stores provided ~47 hours of luminosity
  • Linac personnel repaired a RF station (LRF4)
  • The Cryo System for Pbar's stochastic cooling failed
  • NuMI tripped off on a target LCW indication
  • MI-8 suffered a power glitch that aborted the stash
  • The glitch also caused MI main bus regulation problems

February 21, 2011 - February 23, 2011

  • Three stores provided ~32 hours of luminosity
  • The Antiproton Source worked on their target station LCW system
  • NuMI had problems with their Humidity control system
  • Tevatron personnel conducted beam alignment studies

February 18, 2011 - February 21, 2011

The Integrated Luminosity for the period from 2/14/11 to 2/21/11 was 33.64 inverse picobarns. NuMI reported receiving 7.42E18 protons on target during this same period.

  • Six stores provided ~50.75 hours of luminosity
  • Vacuum personnel fixed a vacuum leak on the Tevatron D17 collimators
  • The Antiproton Source developed with water flow sensor
  • On Monday morning (2/21/11), Pbar personnel accesses and began repairs on two water leaks

February 16, 2011 - February 18, 2011

  • Four stores provided ~30.5 hours of luminosity
  • Pelletron repairs continue
  • NuMI recovered from water skid trip
  • University of Iowa experiment T953 resumed taking beam in Meson
  • Tevatron developed collimator and vacuum problems in sector D

February 14, 2011 - February 16, 2011

  • Three stores provided ~33.5 hours of luminosity
  • Interlock personnel replace a bad MI card
  • Controls personnel replace a bad Pbar card
  • Tevatron cryo system quenches after store 8498 was terminated
  • Booster personnel replaced power supply for low-level RF system
  • Booster personnel repairs termination for Chopper
  • Network Storm hits; Controls personnel located and isolated problem, but the storm affected many pieces of equipment
  • Preaccelerator personnel switched from the H- Source to the I- Source
  • Booster personnel worked on extraction issues
  • D0 required an 8-hour access on 2/16/11 to repair detector

February 11, 2011 - February 14, 2011

The Integrated Luminosity for the period from 2/7/11 to 2/14/11 was 52.78 Pb-1 inverse picobarns. NuMI reported receiving 7.33E18 protons on target during this same period.

  • Five stores provided ~40.5 hours of luminosity
  • MiniBooNE frozen sump discharge line cleared
  • Tevatron lost some luminosity due to earthquake in Chile
  • Store 8488 aborted - no cause found
  • ComEd power glitch caused the loss of store 8492, the stash, the stack, cryo systems, CDF equipment, the Pelletron, and the failure of equipment throughout the accelerators
  • MI safety system tripped off due to a ground fault indication

February 9, 2011 - February 11, 2011

  • Two Tevatron stores provided ~45.5 hours of luminosity
  • Linac personnel worked on LRF1 and LFR2
  • Main Injector ground fault found

February 7, 2011 - February 9, 2011

  • Three stores provided ~29.3 hours of luminosity
  • MiniBooNE suffered three hours of downtime due to system problems
  • Tevatron Personnel conducted orbit kicker tests
  • Interlock personnel conducted Meson Safety system tests
  • Linac personnel repaired a RF station that suffered a power surge
  • On Thursday, 2/10/11, MI will access for 2-3 hours to make RF repairs

February 4, 2011 - February 7, 2011

The Integrated Luminosity for the period from 1/31/11 to 2/7/11 was 43.64 inverse picobarns. NuMI reported receiving 5.14E18 protons on target during this same period.

  • Four stores provided ~37.25 hours of luminosity
  • EE Support replaced VCB relays at Tevatron sectors D3 and A3
  • NuMI vacuum circuit breaker repaired
  • Further Tevatron vacuum circuit breaker inspections halted beam to NuMI and Pbar
  • Meson Cryo Central suffered power outage
  • CHL reported a Cold Box failure - Operations aborted store 8478
  • Booster kicker cable repaired
  • MTest experiment T-978 completes run

February 2, 2011 - February 4, 2011

  • Two stores provided ~33.5 hours of luminosity
  • Kautz Rd substation tripped off
  • Plenty of power problems due to the snow storm
  • Pbar off due to critical device problem
  • Many kicker problems in MI and Recycler
  • Stash lost due to Kautz Rd substation trip, which halted beam to many other areas
  • Store 8467 aborted during end-of-store study
  • NuMI off due to power supply VCB problem
  • Operations restored beam everywhere by the end of the Friday midnight shift, except to NuMI

January 31, 2011 - February 2, 2011

  • Three stores provided ~44.25 hours of luminosity
  • Kicker problems cause loss of antiproton transfers to Recycler
  • FESS finds and later fixes air conditioning problem in MI-31
  • H-Source has problem with extraction
  • Controls fixes communication problem with A0 compressor room
  • MI has safety system tests
  • Linac works on LRF3
  • Power glitches due to storm took down many RF systems, caused the loss of the stash, and dropped MI interlocks
  • Recovering from snowstorm problems

January 28, 2011 - January 31, 2011

The Integrated Luminosity for the period from 1/24/11 to 1/31/11 was 67.3 inverse picobarns. NuMI reported receiving 8.65E18 protons on target during this same period.

  • Six stores provided ~61.5 hours of luminosity
  • The network storm disabled beam to T-978
  • Cryo personnel repaired TEV sector B1 wet engine
  • Controls personnel repaired the NuMI front end that was causing the network storms

January 26, 2011 - January 28, 2011

  • Four stores provided ~28 hours of luminosity
  • Store 8447 quenched due to a separator spark
  • Store 8454 quenched due to kicker pre-fire
  • MTest beamline vacuum problems resolved
  • The accelerator complex suffered some downtime due to a network storm

January 24, 2011 - January 26, 2011

  • Four stores provided ~44.75 hours of luminosity
  • MTest Experiment T-978 off due to beamline vacuum problems
  • NuMI LCW expansion tank repaired
  • Linac switched from the I- Source to the H- Source
  • ARF1 and MRF5 work completed

January 21, 2011 - January 24, 2011

The Integrated Luminosity for the period from 1/17/11 to 1/24/11 was 70.18 inverse picobarns. NuMI reported receiving 8.61E18 protons on target during this same period.

  • Four stores provided ~56.75 hours of luminosity
  • An LCW leak on a Pbar Accumulator magnet required an access to repair
  • Shot setup held up due to noisy readbacks
  • TEV quench during scrape in shot setup
  • Linac personnel will switch from I- Source to H- Source today

January 19, 2011 - January 21, 2011

  • Five stores provided ~36.75 hours of luminosity
  • Store 8430 aborted when a low beta magnet power supply tripped off
  • Problems with a Pbar RF station interrupted stacking for 30 minutes
  • Problems with a MI RF station held off beam to NuMI and Pbar for 30 minutes

January 17, 2011 - January 19, 2011

  • Three stores provided ~45.25 hours of luminosity
  • A network storm hit and caused some downtime

January 14, 2011 - January 17, 2011

The Integrated Luminosity for the period from 1/10/11 to 1/17/11 was 64.13 inverse picobarns. NuMI reported receiving 7.82E18 protons on target during this same period.

  • Five stores provided ~ 45.75 hours of luminosity
  • FESS found serious problems with a CUB chilled water system
  • Store 8423 aborted due to a Tevatron power supply problem

January 12, 2011 - January 14, 2011

  • Four stores provided ~41 hours of luminosity
  • NuMI conducted target and Horn scans
  • Operations rebooted MECAR
  • Store 8417 aborted due to Tevatron RF station trip

January 10, 2011 - January 12, 2011

  • Three stores proved ~44.75 hours of luminosity
  • MI-60 upper bend power supply repaired
  • Short down times caused by Linac power supply trips and Buncher trips

January 7, 2011 - January 10, 2011

The Integrated Luminosity for the period from 1/3/11 to 1/10/11 was 10.79 inverse picobarns. NuMI reported receiving 7.64E18 protons on target during this same period.

  • Three stores provided ~18.5 hours of luminosity
  • The TEV quenched many times over the weekend
  • FTBF T-978 (CALICE) has returned to the MTest beamline
  • A Pbar CAMAC power supply failed and caused the loss of 38mA of beam from the stack

January 5, 2011 - January 7, 2011

  • One store provided ~8.25 hours of luminosity
  • Tevatron recovery continues
  • First Store since Christmas

January 3, 2011 - January 5, 2011

  • No stores due to TEV sector B1 vacuum problems
  • Safety personnel beginning safety system checks
  • TEV sector B1 vacuum problems fixed and cryo system cooldown begins

December 22, 2010 - January 3, 2011

The Integrated Luminosity for the period from 12/22/10 to 1/3/11 was ~53.9 inverse picobarns. NuMI reported receiving 8.81E18 protons from 12/2710 to 1/3/11.

  • Five stores provided ~65.5 hours of luminosity
  • T-932 (Diamond) began taking beam at MTest
  • Store 8388 quenched due to a vacuum problem
  • Cryo system technicians began a warm up of B1 to change out a spool piece
  • NuMI and MiniBooNE continued to take beam as it was available
  • Beam to T-932 stopped due to accesses into the Transfer Hall
  • Machines conducted maintenance and studies


Acclerator Update Archive:
2012 | 2011 | 2010 | 2009 | 2008 | 2007 | 2006 | 2005 | 2004 | 2003

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.


Return to Current Accelerator Update

 
last modified 12/28/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory